OpenStudy (anonymous):

cubic meter of concerte mix contains 390 kg of cement , 160 kg of stone and 110 of sand . what is the total weight of the cubic meter of concrete mix and what part is stone ? the stone part is 660 when u add them all up

OpenStudy (anonymous):

To figure our the totoal weight, just take the sum of each individual weight 390 kg+160 kg+110 kg=660 kg. I don't understand what you mean when you say, "what part is stone?" and I don't understand the last sentence of your question. Why is the stone 660 kg? The first part telling me that the stone part is 160 kg?

OpenStudy (anonymous):

ok we no that but u take them add them all up u get 660/390 and then u do the other part to simp it says 23/3 but thats wrong the right one is 8/33 but how

OpenStudy (anonymous):

Okay. With that being said, I think that last part of your question is what fraction of the total weight of the concrete mix is stone. To figure that out, you just have to divide the weight of the stone (160) by the total weight (660). Simplifying the simplification of that looks like this: 160/660 = 16/66 = 8/33....8/33 is your answer. Side note, you could convert this answer to a fraction that would represent 24.24 % of the concrete mix is stone.

OpenStudy (anonymous):

I mean you could convert to a percentage representation by dividing 8 by 33 and multiplying be 100 to get 24.24%